OpenStudy (anonymous):

The area (in square inches) of a rectangle is given by the polynomial function A(m) = m2 + 7m + 12. If the width of the rectangle is (m + 3) inches, what is the length?

rishavraj (rishavraj):

length*width = area.. u got the area and width , get the length

OpenStudy (anonymous):

how do i do that though with no numbers

rishavraj (rishavraj):

length*(m+3) = m^2+7m+12 \[length = \frac{ m^2+7m+12 }{ m+3 }\] either u can divide the numerator by denominator... of proceed by factorising the numeratior

OpenStudy (mathmale):

"how do i do that though with no numbers?" You'll need to do the necessary algebra, which involves mainly variables, such as x, instead of specific numeric values. See the suggestion of rishavraj , above. Divide m+3 into m^2 + 7m + 12, please. If unsure of what to do, please ask some questions.